OpenStudy (anonymous):

solve the equationto find the value of X 4x+10=22

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

4x+10=22 4x=22-10 4x=12 x=12/4 x=3